Rodney is 55-years-old, stands 6’ tall, weighs 280 pounds and has a 44” waist circumference. He has always had a sedentary job and loves to cook. Both his mother and father are overweight, and his father suffered a heart attack at age 45. During his annual physical exam, Rodney’s doctor diagnosed him with metabolic syndrome and suggested that he may be a candidate for bariatric surgery, as his lab results and family history put him in the high-risk category chronic diseases.

Question 2: Calculate Rodney’s BMI and explain how his waist circumference fits into his risk for chronic disease. Include in your response specific data, facts, examples, key terminology and other information drawn from the textbook and at least one other supplemental source.
Answer #1

BMI=703×Weight (lbs)÷Height(inch²)

Here weight=280 lbs

Height=6 (72inch) as 1'=12"

Therefore

=703×280÷(72)²

=196,840÷(72)²

=196,840÷5184

=37.9 or 38

The BMI is 38kg/m²

A BMI above 35kg/m² makes an individual very high risk patient because of obesity. They are classified under morbid obesity. They often needs immediate usually surgical treatment to maintain their weight and prevent complications

The males if they have a waist circumference above 40inches are considered to be very obese and are very high risk individual for having cardiovascular issues and type 2 diabetes.

In this patient the BMI and the waist circumference both are abnormal or very much elevated qualifying him to be a candidate fie bariatric surgery.